OpenStudy (anonymous):

A quadratic equation is shown below: 3x2 − 16x + 2 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 9x2 + 3x − 2 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. (5 points)

RhondaSommer (rhondasommer):

Part A) The discriminant of 3x2 − 16x + 2 is 232. what does that tell us?

OpenStudy (anonymous):

that its positive?

OpenStudy (anonymous):

im very confused on how to do this

RhondaSommer (rhondasommer):

okay so the radicand tells us how many soulutions will have

RhondaSommer (rhondasommer):

we'll*

OpenStudy (anonymous):

2

RhondaSommer (rhondasommer):

yes. That tells us there are 2 real solutions there is your answer to part a

RhondaSommer (rhondasommer):

Part B) solve by factoring 9x2 + 3x − 2 = 0 (3x+2)(3x-1)=0 3x+2 = 0 or 3x-1 = 0

OpenStudy (anonymous):

but for part a it says to show work i didnt really do any work

RhondaSommer (rhondasommer):

one second

OpenStudy (anonymous):

???

RhondaSommer (rhondasommer):

okay back

RhondaSommer (rhondasommer):

\[ax^2+bx+c\] \[3x^2-16x+2\] what is a, b, and c

OpenStudy (anonymous):

a= 3 b=-16 and c= 2

RhondaSommer (rhondasommer):

we will use discrimanant b^2-4ac which equals what?

RhondaSommer (rhondasommer):

plug in the letters for the numbers

OpenStudy (anonymous):

would it be 256-24?

RhondaSommer (rhondasommer):

the value under the square root (b²-4ac) is called Discriminant (what i just gave you) and its value tells us about the nature of the roots of the equation. If D = 0, the roots are real & equal if D > 0, the roots are real & different if D< 0, the roots are imaginary & different (yes you are correct)

OpenStudy (anonymous):

ok thats for part "b" right?

RhondaSommer (rhondasommer):

not really thajust for when you get your answers to part b

RhondaSommer (rhondasommer):

Part A) The discriminant of 3x2 − 16x + 2 is 232. That tells us there are 2 real solutions Part B) solve by factoring 9x2 + 3x − 2 = 0 (3x+2)(3x-1)=0 equal each to 0 3x+2 = 0 or 3x-1 = 0 -2/3=x 1/3=x

RhondaSommer (rhondasommer):

hello?

OpenStudy (anonymous):

hi sorry my computer died

OpenStudy (anonymous):

@RhondaSommer

RhondaSommer (rhondasommer):

back

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

back to the problem

OpenStudy (anonymous):

@rhondasommer

RhondaSommer (rhondasommer):

yes sorry i was busy

OpenStudy (anonymous):

well i did the question by myself i probably got it wrong but whatever thanks though. @RhondaSommer

RhondaSommer (rhondasommer):

I legit gave you the answers \_(-.-)_/